Question 6 (1 point) Saved Suppose that one-way commute times in a particular city are normally distributed with a mean of 18.08 minutes and a standard deviation of 2.661 minutes. Would it be unusual for a commute time to be above 21.6 minutes? 0 1) The value is borderline unusual. 2) We do not have enough information to determine if the value is unusual 3) The value is not unusual 4) It is impossible for this value to occur with this distribution of data. 5) The value is unusual.
Answer #1

Solution :

Given that,

mean = \mu = 18.08

standard deviation = \sigma = 2.661

P ( x > 21.6 )

= 1 - P (x < 21.6 )

= 1 - P ( x -  \mu/ \sigma) < ( 21.6 - 18.08 / 2.661)

= 1 - P ( z < 3.52 / 2.661 )

= 1 - P ( z < 1.32)

Using z table

= 1 - 0.9066

= 0.0934

Probability = 0.0934

The value is not unusual.